OpenStudy (anonymous):

Find all the quartiles for the following data set. 32, 54, 32, 33, 59, 38, 46, 44

OpenStudy (angelic):

put it in order

OpenStudy (angelic):

Arrange the data in a numerical order. 32 32 33 38 44 46 54 59 The median or the second quartile: (38+44)/2 = 41. 32 32 33 38 | 44 46 54 59 The lower or first quartile: (32+33)/2 = 32.5 32 32 | 33 38 | 44 46 54 59 The upper or third quartile: (46+54)/2 = 50 32 32 | 33 38 | 44 46 | 54 59